OpenStudy (anonymous):

F=(1/a)-(s^2 / a^2)*F Trying to solve for F

OpenStudy (anonymous):

So far I have gotten after a bit of algebra \[F = \frac{ 1 }{ a(1+\frac{ s^2 }{ a^2 }) } \] I mus be getting sleepy however, as I still can't see how this becomes \[F = \frac{ a }{ s^2 + a^2 }\]

OpenStudy (anonymous):

Nevermind. Figured it out myself. Simplify, Then multiply top and bottom by a